# DELIFT: Data Efficient Language model Instruction Fine-Tuning

### v2

@@ A.1 General Observations
A.2 Detailed Analysis of Methods
A.2.1 Initial vs. Random Selection
@@ A.2 Detailed Analysis of Methods
A.2.2 SelectIT and LESS Methods
@@ A.2.3 DELIFT Variants
A.3 Impact on Different Fine-Tuning Scenarios
@@
Appendix B Prometheus Rubric
@@ B.1 Usage Notes
Appendix C LLM-as-Judges Scores
In Tables 8 and 9, we show the distribution of Prometheus scores on one particular setting: Use Case 1, MixInstruct training and MixInstruct validation sets on the Qwen2-72B-Instruct model. These figures make clear that the average LGA scores computed in Tables 1-6 are true averages of a distribution of scores, not averages of a combination of just 1’s and 5’s.
@@ Appendix C LLM-as-Judges Scores
C.1 Interpretation of Score Distributions
C.1.1 Overall Trends
@@
C.1.2 Method-Specific Observations
@@
C.1.3 Comparison with Full Data
@@ Appendix D Limitations
Appendix E Future Work
@@
Appendix F Code and Data Availability
To facilitate reproducibility and further research, we will make the DELIFT implementation and the datasets used in our experiments publicly available upon publication. Interested researchers can access these resources through the following repository: https://anonymous.4open.science/r/optimizing-data-selection-0CD0.
Appendix G Hyperparameter Settings

### v3

@@ B.1 General Observations
B.2 Detailed Analysis of Methods
B.2.1 Initial vs. Random Selection
@@ B.2 Detailed Analysis of Methods
B.2.2 SelectIT and LESS Methods
@@ B.2.3 DELIFT Variants
B.3 Impact on Different Fine-Tuning Scenarios
@@
Appendix C Prometheus Rubric
@@ C.1 Usage Notes
Appendix D LLM-as-Judges Scores
In Tables 11 and 12, we show the distribution of Prometheus scores on one particular setting: Use Case 1, MixInstruct training and MixInstruct validation sets on the Qwen2-72B-Instruct model. These figures make clear that the average LGA scores computed in Tables 1-8 are true averages of a distribution of scores, not averages of a combination of just 1’s and 5’s.
@@ Appendix D LLM-as-Judges Scores
D.1 Interpretation of Score Distributions
D.1.1 Overall Trends
@@
D.1.2 Method-Specific Observations
@@
D.1.3 Comparison with Full Data
@@ Appendix E Limitations
Appendix F Future Work
@@
Appendix G Code and Data Availability
To facilitate reproducibility and further research, we will make the DELIFT implementation and the datasets used in our experiments publicly available. Interested researchers can access these resources through the following repository: https://anonymous.4open.science/r/optimizing-data-selection-0CD0.
Appendix H Hyperparameter Settings